OpenStudy (anonymous):

What is the solution to the equation log(2) x + log(2) 4 = 3 ?

Parth (parthkohli):

A good start is knowing what \(\log _24 \) equals.

OpenStudy (anonymous):

I know i have to re-write the equation but i don't know what to change it to to make it... addable.

Parth (parthkohli):

Never worry about that. Can you tell me what \(\log_2 4\) is equal to?

Parth (parthkohli):

You don't need a calculator to calculate \(\log_24\). What power of 2 is 4?

OpenStudy (anonymous):

2

Parth (parthkohli):

There you go. So \(\log_24 = 2\). Now we have something like \(\log_2x + 2 = 3\).

Parth (parthkohli):

Subtract 3 from both sides to start with.

Parth (parthkohli):

\(\log_2x = 3 - 2\). My mistake - subtract 2 from both sides.

OpenStudy (anonymous):

so log(2) x=1 which means x is .10? or is that wrong?

Parth (parthkohli):

\(\log_2x = 1\) This just means that 2 to some power is 1.

Parth (parthkohli):

What power of \(2\) is \(1\)?

OpenStudy (anonymous):

Blah .5.

OpenStudy (anonymous):

so x is equal to .5

Parth (parthkohli):

How is \(2^{0.5} = 1?\)

OpenStudy (lgbasallote):

careful

OpenStudy (lgbasallote):

\[\Large \log_2 x = 1 \implies 2^1 = x\]

Parth (parthkohli):

Hmm.

Parth (parthkohli):

What lgbasallote said.

Parth (parthkohli):

What is \(2^1\)?

OpenStudy (lgbasallote):

oh man..long have i waited to say that \m/

Parth (parthkohli):

I see.

OpenStudy (anonymous):

isn't 2^1=2?

Parth (parthkohli):

Yep

Parth (parthkohli):

So that's it. \(x = 2\).

OpenStudy (anonymous):

But then dosn't log(2)2+log(2)4 not equal 3? it would equal 4.

Parth (parthkohli):

How so?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \log_22 = 1, \log 2_4 = 2 \Longrightarrow 2 + 1 = 3}\)

OpenStudy (anonymous):

oh i see now! I'm a moron sorry! thank you for your help :)

Parth (parthkohli):

Heh. You're welcome!